Find (p+q)÷(p-q),if:
p=1/4,q=3/2 what is the answer of this question tell me about it please 

Grade:8

2 Answers

Armaan Arora
3 years ago
p=1/4
q=3/2
p+q= 1/4+3/2 = 1+6/4 = 7/4
p-q= ¼ – 3/2  = -5/2
p+q/p-q= 7/4/-5/2 = 7/4*2/-5 = -7/10
hope you understand
Sathwik reddy
3 years ago
for  a smart method use componendo and dividendo method
first find
p/q = 1/6
now apply componendo dividendo
we get p+q/p-q =1+6/1-6
finally the ans is 
p+q/p-q =-7/5